OpenStudy (anonymous):

A rectangular field is 39m by 30m. Find its area in m^2 and in ha. I know that the area in m^2 is 1170m^2 but how do I find the ha

jimthompson5910 (jim_thompson5910):

Hint: 1 hectare = 10,000 square meters

jimthompson5910 (jim_thompson5910):

Note: ha is the abbreviation of a hectare

OpenStudy (anonymous):

yes I understand that but how do I find the area in ha?

jimthompson5910 (jim_thompson5910):

Multiply 1170 by 1/(10000) to convert to hectares, so... 1170*(1/10000) = 1170/10000 = 0.117 Therefore, 1170 square meters = 0.117 hectares

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome